Suppose that a certain expression is represented by f(x) and that f(3) = -5. What is one point of the graph of f ?
Answer by DrBeeee(684)

When we have an equation, for example,
(1) y = 2x - 11
we are specifically naming the function y. To be more correct, we should start with
(2) f(x) = 2x - 11, then name it as
(3) y = f(x) or
(4) y = 2x - 11.
In actuality, the name of the function, 2x-11, is plain "f". I'm sure you've seen the use of g(x). For example
(5) g(x) = -3x + 2, this function is called (named) "g".
The names f and g specify what is called the topology of the mathematical expression, (2) and (5) respectively. In Algebra I we tend to use only the equation portion to solve problems. In advanced math we use functions.
So much for the tutorial, let's get to your question.
You ask for a point on the graph of f(x) vs x. The reason for my initial comments above, was to see that you are asking for a point on the graph of y vs x. I'm sure you have drawn some by now!
Let's assume that (1) is the function you are using. When x =3
(6) y = f(x) or
(7) y = f(3) or
(8) y = 2*3 - 11 or
(9) y = 6 - 11 or
(10) y = -5
So you would "draw" the point (x,y) = (3,-5), right?
Answer: the point is (x,f(x)) = (3,-5).
PS I invented function of (1) to show what you already know to answer the question.